# SELFI: Autonomous Self-Improvement with Reinforcement Learning for Social Navigation

### v1

@@ Abstract
I Introduction
Reinforcement learning (RL) provides an appealing algorithmic approach for autonomously improving robotic policies in unpredictable and complex real-world settings [29, 42, 27, 11]. For example, in the indoor navigation scenario depicted in Fig. 1, the robot needs to not only avoid obstacles, but also deal with unpredictable and hard-to-model situations, like the interaction with the pedestrian. Model-based control methods can struggle with the unpredictable elements in the scene, and RL in principle provides an appealing alternative: learn directly from real-world experience, sidestepping the need for highly accurate modeling. However, directly performing end-to-end RL from scratch in the real world can be difficult: discovering a high-quality policy may require a large number of trials and encounter catastrophic failures during the training process [21, 25, 26]. This is especially problematic when it is not possible to provide external instrumentation that avoids catastrophic failures — for example, with human interactions where failures might be inconvenient or even dangerous.
@@ I Introduction
We illustrate this design in Fig. 1, and we instantiate our system in the context of social navigation: the problem of navigating an indoor space while avoiding undesirable behavior around pedestrians, such as interruptions, collisions, or penetration of their personal space. This problem domain is a good fit for validating our framework because model-based policies can be derived from geometric models of the world and rough predictive models of pedestrians, but these policies can be significantly improved through model-free trial-and-error, both because the pedestrian models might be inaccurate, and because the robot can adapt directly to the behavior of the pedestrians in a specific downstream deployment environment. In this setting, our model-based policies are derived from the previously proposed SACSoN method [18], which constructs policies by optimizing a trajectory value using a 3D reconstruction of training environments and predictive models of humans. This model-based procedure also provides trajectory value estimates that can bootstrap the model-free RL critic. During the real-world model-free RL phase, we improve the model-based policy by learning a residual value critic and applying actor-critic methods as described above.
@@
The main contributions of this paper is to propose a framework, SELFI, that takes advantage of the best aspects of online RL and offline model-based learning. Specifically, SELFI uses online Q-learning to fine-tune a control policy trained with offline model-based learning. SELFI rapidly improves the performance of a pre-trained policy in the target environment without needing significant human intervention during online learning. In our evaluations, SELFI improves the performace of the pre-trained policy in multiple vision-based navigation tasks, greatly outperforming policies trained purely offline as well as standard end-to-end model-free offline-to-online RL finetuning methods. Within only two hours of fine-tuning, our policy learns complex robotic behaviors, e.g., pre-emptive behavior for navigating around pedestrians, collision avoidance for unseen small and transparent obstacles, and preferences for smooth and easily traversable surfaces. The robotic behaviors learned by SELFI are shown in our project page.
II Related Work
@@
There are various data-driven methods for adapting control policies to their real-world environments through interactions. In classical control settings, adaptive control [3] and iterative learning control [5] are widely studied methods. In learning-based settings, DAgger [39] is a general framework that iteratively trains a control policy with expert labeled demonstrations. Additionally, model-based learning can fine-tune control policies by utilizing differentiable dynamic forward models to define an objective function with visual foresight [35, 15, 16], reward prediction [12], and state prediction [43, 17, 18]. This type of learning can be combined with an optimization algorithm to generate action commands online, just as in model predictive control [7, 6, 13, 4]. However, for any model-based approach, the performance of the learned control policy is limited to the accuracy of the model and the quality of the dataset. Model-based RL, which learns a model from interactions with the real world, is subject to similar limitations [30, 36, 20].
@@
Our proposed approach is closely related to Residual RL [19], which decomposes the policy’s output into a solution from an existing controller and the actions from a residual policy trained with model-free RL. By leveraging the existing controller, residual RL stabilizes the robot’s behavior during the early stages of online learning and learns the target behavior via the flexibility of model-free RL. However, the confusion between the existing and learned control policies due to composing them in the action space restricts the performance of residual RL. Different from these previous works, our proposed method, SELFI, seamlessly composes model-based learning and model-free RL in the objective space. SELFI is a flexible and stable method for online fine-tuning with model-free RL because it incorporates the objective used in offline model-based learning in the Q-values it learns online.
III Combining Model-based Control with Online Model-Free RL
III-A Preliminaries
@@
Model-free RL: In RL, we want to maximize the expected sum of discounted rewards. $Q$-learning algorithms [28, 11] solve this task by learning a function approximating
$Q^{\pi_{\theta}}(s,a)=\operatorname*{\mathop{\mathbb{E}}}_{\{s_{t},a_{t}\}\sim\mathcal{M}^{\pi}}\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t}),$ (1)
@@
where trajectories with states $s_{t}$ and actions $a_{t}$ are sampled from the closed-loop dynamics of $\mathcal{M}$ under the policy $\pi_{\theta}$. $\gamma$ indicates the discount factor for future rewards $r$. Here, mathematical symbols without a subscript representing time show the current time state and action, e.g., $s=s_{0}$, $a=a_{0}$. The $Q$-function for the optimal policy obeys the Bellman equation, and it can be trained to minimize the TD error $\delta$:
$\delta=r(s,a)+\gamma\max_{a^{\prime}}Q^{\pi_{\theta}}(s^{\prime},a^{\prime})-Q^{\pi_{\theta}}(s,a).$ (2)
where $s^{\prime}$ and $a^{\prime}$ indicate the next step $s$ and $a$, respectively. In the actor-critic setting, we learn both an approximation for the action-value function $Q^{\pi_{\theta}}(s,a)$ and for the policy that maximizes the action-value function $Q(s,a)$ as $\pi_{\theta}(s)$. This enables tractable optimization over large action spaces where the $\max$ in the Bellman equation cannot be efficiently computed.
Model-based learning: In the model-based learning setting, we optimize an objective over open-loop sequences of virtual actions $\tau=\{\hat{a}_{t}\}_{t=0\ldots H-1}$ using an approximate dynamics model and reward estimate:
$\operatorname*{arg\,max}_{\tau=\{\hat{a}_{t}\}}\left[J(s,\tau):=\sum_{t=0}^{H-1}\gamma^{t}\hat{r}(\hat{s}_{t},\hat{a}_{t})\right]\textrm{ s.t. }\hat{s}_{t+1}=\hat{f}(\hat{s}_{t},\hat{a}_{t}),$ (3)
@@
where $\hat{s}_{t}$ is the predicted state under the approximate dynamics described by $\hat{f}$, which can be either learned or given. $\hat{a}_{t}$ is the $t-$th virtual action, which is estimated at the current time state. Note that $\hat{s}_{0}=s$ and $\hat{a}_{0}=a$ in Eqn 3. We assume that the dynamics $\hat{s}_{t+1}=\hat{f}(\hat{s}_{t},\hat{a}_{t})$ and rewards $\hat{r}(\hat{s}_{t},\hat{a}_{t})$ are easy to compute and differentiable, allowing us to directly optimize the sequence of actions online with gradient-based trajectory optimization [7, 13]. Instead of optimizing this objective online (analogous to traditional nonlinear model-predictive control methods), model-based policy learning amortizes this optimization by learning the parameters $\theta$ of a control policy $\tau=\pi_{\theta}(s)$ to optimize this objective offline. This distills the optimization problem into an offline base policy $\pi_{\theta}$, represented as a neural network mapping observations to sequences of actions $\tau$. At runtime, actions are then sampled from the learned policy $\pi_{\theta}$. A number of prior methods have proposed similar methods for model-based policy learning [35, 12, 15, 17].
III-B SELFI learning architecture
In SELFI, we combine the strengths of model-based learning with the benefits of model-free learning to enable finetuning in the real world. We wish to decompose the critic value into a model-based component $J$ and a learned residual objective $\bar{Q}$:
$Q(s,\tau)=J(s,\tau)+\bar{Q}(s,a)$ (4)
@@ III-B SELFI learning architecture
where $a$ is the first action $\hat{a}_{0}$ in the sequence of virtual actions $\tau=\{\hat{a}_{0},\ldots,\hat{a}_{H-1}\}$. We train $\bar{Q}(s,a)$, which corresponds to the part of the overall objective that is not considered by the model-based learning, and fine-tune the pre-trained control policy $\pi_{\theta}(s)$ to maximize the combined Q-function $Q(s,\tau)$. Accordingly, we first assume $r(s_{t},a_{t})$ in timesteps $t=0\ldots H-1$ has the form:
$r(s_{t},a_{t})=\hat{r}(\hat{s}_{t},\hat{a}_{t})+\bar{r}(s_{t},a_{t}),$ (5)
where $\hat{r}(\hat{s}_{t},\hat{a}_{t})$ is the model-based reward and $\bar{r}(s_{t},a_{t})$ is the unmodeled residual reward. Hence, we have:
$Q(s,\tau)$ $=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$ (6)
$=\underbrace{\sum_{t=0}^{H-1}\gamma^{t}\hat{r}(s_{t},a_{t})}_{J(s,\tau)}+\underbrace{\sum_{t=0}^{H-1}\gamma^{t}\bar{r}(s_{t},a_{t})+\sum_{t=H}^{\infty}\gamma^{t}r(s_{t},a_{t})}_{\bar{Q}(s,a)}$
The model-free $\bar{Q}$ includes both:
  1. 1.
    the reward terms that cannot be directly modeled; and
  2. 2.
    the value of the long-horizon returns that are ignored by the limited-horizon model-based learning.
As usual, $Q(s,\tau)$ should obey the Bellman equation:
$Q(s,\tau)=r(s,a)+\gamma Q(s^{\prime},\pi_{\theta}(s^{\prime})),$ (7)
and the critic for $\bar{Q}(s,a)$ in Eqn 4 can be trained to minimize the TD error $\tilde{\delta}$:
@@
In practice we use a delayed copy of $\bar{Q}$ as a target network to compute target values [10]. During online learning, we calculate the gradient of $\pi_{\theta}$ by back-propagation to maximize the hybrid objective $Q(s,\tau)$ in Eqn 4. By including the model-based objective term $J(s,\tau)$ in $Q(s,\tau)$, the initial estimate of $Q(s,\tau)$ already provides reasonable values with a suitably initialized $\bar{Q}(s,a)$ (e.g., with small initial weights), which significantly stabilizes performance early in training. Then, $\bar{Q}(s,a)$ can be further trained with online interactions and enforce to learn the target robotic behaviors.
Figure 2: System diagram of SELFI. SELFI, implemented in the workstation, trains the actors by maximizing the proposed hybrid objectives and sends the actor parameters to the robot controller. The robot actuates using the trained actors and sends new data to the workstation.
III-C SELFI implementation
@@
As our underlying RL algorithm, we use a variant of twin delayed deep deterministic policy gradients (TD3) [10], which is a sample-efficient and stable algorithm for training deterministic control policies. Fig. 2 shows the system diagram of SELFI for real-world learning. Similar to [46], policy training is implemented on an off-board workstation while the robot’s onboard computer is used for running inference.
@@
On the workstation, we train with batch data from both the offline and online dataset to avoid overfitting to the online data. Half of the batch data is chosen from the offline dataset, and the other half is from online data, which is collected by the robot. We update the actor once for every two updates to the critic, maximizing $Q(s,\tau)$ from Eqn 4 with respect to the policy.
@@
On the robot, we calculate the learned policy to obtain the sequence of actions $\tau=\{\hat{a}_{t}\}_{t=0\ldots H-1}$ and execute the first action $\hat{a_{0}}$, adding Gaussian noise $\epsilon$ to encourage exploration. Then, the robot sends the collected data to the workstation, where it is stored in the replay buffer. Periodically, the updated actor weights are sent from the workstation to the robot so it can use the latest policy when collecting data.
IV SELFI System Setup
@@
We evaluate SELFI on for a vision-based social navigation task, where a robot must navigate an indoor environment with pedestrians. We employ SACSoN [18] as the offline model-based objective. In the online phase, SELFI fine-tunes the pre-trained policy to learn socially-compliant behavior including:
  1. 1.
    pre-emptive avoidance of oncoming pedestrians
  2. 2.
    collision avoidance for the small or transparent objects
  3. 3.
    avoiding travel on uneven floor surfaces.
These behaviors are difficult to learn purely from offline model-based learning due to the modeling errors and insufficient information in the offline dataset. In this section, we describe the implementation of SELFI on top of SACSoN and our hardware setup for robotic navigation.
IV-A Offline Learning with SACSoN

### v2

# SELFI: Autonomous Self-Improvement with RL for Vision-Based Navigation around People

@@ Abstract
Keywords: online reinforcement learning, vision-based navigation
1 Introduction
Reinforcement learning (RL) provides an appealing algorithmic approach for autonomously improving robotic policies in unpredictable and complex real-world settings [1, 2, 3, 4]. For example, in the indoor navigation scenario depicted in Fig. 1, the robot needs to not only avoid obstacles, but also deal with unpredictable and hard-to-model situations, like the interaction with the pedestrian. Model-based control methods can struggle with the unpredictable elements in the scene [5, 6, 7], and RL in principle provides an appealing alternative: learn directly from real-world experience, sidestepping the need for highly accurate modeling. However, directly performing end-to-end RL from scratch in the real world can be difficult: discovering a high-quality policy may require a large number of trials and encounter catastrophic failures during the training process [8, 9, 10]. This is especially problematic when it is not possible to provide external instrumentation that avoids catastrophic failures — for example, with human interactions where failures might be inconvenient or even dangerous.
@@ 1 Introduction
We illustrate this design in Fig. 1, and we instantiate our system in the context of social navigation: the problem of navigating an indoor space while avoiding undesirable behavior around pedestrians, such as interruptions, collisions, or invasion of their intimate distance. This problem domain is a good fit for validating our framework because model-based policies can be derived from geometric models of the world and rough predictive models of pedestrians, but these policies can be significantly improved through model-free trial-and-error, both because the pedestrian models might be inaccurate, and because the robot can adapt directly to the behavior of the pedestrians in a specific downstream deployment environment. In this setting, our model-based policies are derived from the previously proposed SACSoN [5], which constructs policies by optimizing a trajectory value using a 3D reconstruction of training environments and predictive models of humans. This model-based procedure also provides trajectory value estimates that can bootstrap the model-free RL critic. During the real-world model-free RL phase, we improve the model-based policy by learning a residual value critic and applying actor-critic methods as described above.
@@
The main contributions of this paper is to propose a framework, SELFI, that takes advantage of the best aspects of online RL and offline model-based learning. Specifically, SELFI uses online Q-learning to fine-tune a control policy trained with offline model-based learning. SELFI rapidly improves the performance of a pre-trained policy in the target environment without needing significant human intervention during online learning. In our evaluations, SELFI improves the performace of the pre-trained policy in multiple vision-based navigation tasks, greatly outperforming policies trained purely offline as well as standard end-to-end model-free offline-to-online RL finetuning methods. Within only two hours of fine-tuning, our policy learns complex robotic behaviors, e.g., pre-emptive behavior for navigating around pedestrians, collision avoidance for unseen small and transparent obstacles, and preferences for smooth and easily traversable surfaces. The robotic behaviors learned by SELFI are shown in the supplemental videos.
2 Related Work
We review the related learning methods as well as the navigation methods on the evaluation task.
@@ 2 Related Work
Online learning: There are various data-driven methods for adapting control policies to their real-world environments through interactions. In learning-based settings, DAgger [19] is a general framework that iteratively trains a control policy with expert labeled demonstrations. Additionally, model-based learning can fine-tune control policies by utilizing differentiable dynamic forward models to define an objective function with visual foresight [20, 21, 22], reward prediction [23], and state prediction [24, 25, 5]. This type of learning can be combined with an optimization algorithm to generate action commands online, just as in model predictive control [26, 27, 28, 29]. However, for any model-based approach, the performance of the learned control policy is limited to the accuracy of the model and the quality of the dataset. Model-based RL, which learns a model from interactions with the real world, is subject to similar limitations [30, 31, 32].
@@
Our proposed approach is closely related to Residual RL [14], which decomposes the policy’s output into a solution from an existing controller and the actions from a residual policy trained with model-free RL. By leveraging the existing controller, residual RL stabilizes the robot’s behavior during the early stages of online learning and learns the target behavior via the flexibility of model-free RL. However, the confusion between the existing and learned control policies due to composing them in the action space restricts the performance of residual RL. Different from these previous works, our proposed method, SELFI, seamlessly composes model-based learning and model-free RL in the objective space. SELFI is a flexible and stable method for online fine-tuning with model-free RL because it incorporates the objective used in offline model-based learning in the learned Q-values.
Social navigation: Social navigation has been extensively explored in [41, 42, 43]. Model-based approaches relying on dynamic pedestrian models have been utilized for behavior modeling [44, 7, 6, 45, 46]. Many existing techniques determine the robot’s actions based on predicted pedestrian behavior [47, 48, 49, 50, 51, 52, 53]. Furthermore, social navigation has also been explored using model-free learning approaches, such as reinforcement learning [54, 55, 56, 57, 58, 59, 60]. Unlike vision-based navigation [61, 62, 63, 21, 25] as our evaluation task, which only uses RGB camera observations, these approaches rely on detected pedestrian poses and/or multiple metric sensors such as LiDARs and/or depth cameras [64, 65].
The most related work, SACSoN [5], utilizes the model-based learning architecture to learn a socially unobtrusive policy in vision-based navigation setting. However, even simple socially compliant behaviors are still challenging to learn due to the modeling errors. Hence, we apply our proposed method, SELFI, to fine-tune the pre-trained SACSoN policy to improve its performance. We conduct comparisons with baseline methods as well as other online fine-tuning methods in the same task setting, vision-based navigation as our method.
3 Combining Model-based Control with Online Model-Free RL
3.1 Preliminaries
@@
Model-free RL: In RL, we want to maximize the expected sum of discounted rewards. $Q$-learning algorithms [66, 4] solve this task by learning a function approximating $Q^{\pi_{\theta}}(s,a)=\operatorname*{\mathop{\mathbb{E}}}_{\{s_{t},a_{t}\}\sim\mathcal{M}^{\pi}}\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$ where trajectories with states $s_{t}$ and actions $a_{t}$ are sampled from the closed-loop dynamics of $\mathcal{M}$ under the policy $\pi_{\theta}$. $\gamma$ indicates the discount factor for future rewards $r$. Here, mathematical symbols without a subscript representing time show the current time state and action, e.g., $s=s_{0}$, $a=a_{0}$. The $Q$-function for the optimal policy obeys the Bellman equation, and it can be trained to minimize the TD error $\delta=r(s,a)+\gamma\max_{a^{\prime}}Q^{\pi_{\theta}}(s^{\prime},a^{\prime})-Q^{\pi_{\theta}}(s,a)$ where $s^{\prime}$ and $a^{\prime}$ indicate the next step $s$ and $a$, respectively. In the actor-critic setting, we learn both an approximation for the action-value function $Q^{\pi_{\theta}}(s,a)$ and for the policy that maximizes the action-value function $Q(s,a)$ as $\pi_{\theta}(s)$. This enables tractable optimization over large action spaces where the $\max$ in the Bellman equation cannot be efficiently computed.
@@
Model-based learning: In the model-based learning setting, we optimize an objective over open-loop sequences of virtual actions $\tau=\{\hat{a}_{t}\}_{t=0\ldots H-1}$ using an approximate dynamics model and reward estimate as $\operatorname*{arg\,max}_{\tau=\{\hat{a}_{t}\}}\left[J(s,\tau):=\sum_{t=0}^{H-1}\gamma^{t}\hat{r}(\hat{s}_{t},\hat{a}_{t})\right]\textrm{ s.t. }\hat{s}_{t+1}=\hat{f}(\hat{s}_{t},\hat{a}_{t})$ where $\hat{s}_{t}$ is the predicted state under the approximate dynamics described by $\hat{f}$, which can be either learned or given. $\hat{a}_{t}$ is the $t-$th virtual action, which is estimated at the current time state. Note that $\hat{s}_{0}=s$ and $\hat{a}_{0}=a$ in $J(s,\tau)$. We assume that the dynamics $\hat{s}_{t+1}=\hat{f}(\hat{s}_{t},\hat{a}_{t})$ and rewards $\hat{r}(\hat{s}_{t},\hat{a}_{t})$ are easy to compute and differentiable, allowing us to directly optimize the sequence of actions online with gradient-based trajectory optimization [26, 28]. Instead of optimizing this objective online (analogous to traditional nonlinear model-predictive control methods), model-based learning amortizes this optimization by learning the parameters $\theta$ of a control policy $\tau=\pi_{\theta}(s)$ to optimize this objective offline. This distills the optimization problem into an offline base policy $\pi_{\theta}$, represented as a neural network mapping observations to sequences of actions $\tau$. At runtime, actions are then sampled from the learned policy $\pi_{\theta}$. A number of prior methods have proposed similar methods for model-based learning [20, 23, 21, 25].
3.2 SELFI learning architecture
@@
In SELFI, we combine the strengths of model-based learning with the strengths of model-free learning to enable finetuning in the real world. We wish to decompose the critic value $Q$ into a model-based objective $J$ and a learned residual objective $\bar{Q}$ as $Q(s,\tau)=J(s,\tau)+\bar{Q}(s,a)$ where $a$ is the first action $\hat{a}_{0}$ in the sequence of virtual actions $\tau=\{\hat{a}_{0},\ldots,\hat{a}_{H-1}\}$. We train $\bar{Q}(s,a)$, which corresponds to the part of the overall objective that is not considered by model-based learning. We fine-tune the pre-trained control policy $\pi_{\theta}(s)$ to maximize the combined Q-function $Q(s,\tau)$. Accordingly, we first assume $r(s_{t},a_{t})$ in timesteps $t=0\ldots H-1$ has the form, $r(s_{t},a_{t})=\hat{r}(\hat{s}_{t},\hat{a}_{t})+\bar{r}(s_{t},a_{t})$ where $\hat{r}(\hat{s}_{t},\hat{a}_{t})$ is the model-based reward and $\bar{r}(s_{t},a_{t})$ is the unmodeled residual reward. Hence, we have:
$Q(s,\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})=\underbrace{\sum_{t=0}^{H-1}\gamma^{t}\hat{r}(s_{t},a_{t})}_{J(s,\tau)}+\underbrace{\sum_{t=0}^{H-1}\gamma^{t}\bar{r}(s_{t},a_{t})+\sum_{t=H}^{\infty}\gamma^{t}r(s_{t},a_{t})}_{\bar{Q}(s,a)}.$ (1)
The model-free $\bar{Q}$ includes i) the reward terms that cannot be directly modeled and ii) the value of the long-horizon returns that are ignored by the limited-horizon model-based learning. As usual, $Q(s,\tau)$ should obey the Bellman equation $Q(s,\tau)=r(s,a)+\gamma Q(s^{\prime},\pi_{\theta}(s^{\prime}))$ and the critic for $\bar{Q}(s,a)$ can be trained to minimize the TD error $\tilde{\delta}$:
@@
In practice we use a delayed copy of $\bar{Q}$ as a target network to compute target values [18]. During online learning, we calculate the gradient of $\pi_{\theta}$ by back-propagation to maximize the hybrid objective $Q(s,\tau)$. By including the model-based objective term $J(s,\tau)$ in $Q(s,\tau)$, the initial estimate of $Q(s,\tau)$ already provides reasonable values with a suitably initialized $\bar{Q}(s,a)$ (e.g., with small initial weights), which significantly stabilizes performance early in training. Then, $\bar{Q}(s,a)$ can be further trained with online interactions to learn the target robotic behaviors.
3.3 SELFI implementation
@@
As our underlying RL algorithm, we use a variant of twin delayed deep deterministic policy gradients (TD3) [18], which is a sample-efficient and stable algorithm for training deterministic control policies. Fig. 1 shows the system overview of SELFI for real-world learning. Similar to [13], policy training is implemented on an workstation while the robot’s onboard computer is used for inference.
@@
On the workstation, we train with batch data from both the offline and online dataset to avoid overfitting to the online data. Half of the batch data is chosen from the offline dataset, and the other half is from online data, which is collected by the robot. We update the actor once for every two updates to the critic, maximizing $Q(s,\tau)=J(s,\tau)+\bar{Q}(s,a)$ with respect to the policy.
@@
On the robot, we calculate the learned policy to obtain the sequence of actions $\tau=\{\hat{a}_{t}\}_{t=0\ldots H-1}$ and execute the first action $\hat{a_{0}}$, adding Gaussian noise $\epsilon$ to encourage exploration. Then, the robot sends the collected data to the workstation, where it is stored in the replay buffer. Periodically, the updated actor weights are sent from the workstation to the robot so it can use the latest policy.
4 SELFI System Setup
@@
We evaluate SELFI on for a vision-based social navigation task, where a robot navigates an indoor environment with pedestrians. We employ SACSoN [5] as the offline model-based objective. In the online phase, SELFI fine-tunes the pre-trained policy to learn socially-compliant behavior including 1) pre-emptive avoidance of oncoming pedestrians, 2) collision avoidance for the small or transparent objects, and 3) avoiding travel on uneven floor surfaces. These behaviors are difficult to learn purely from offline model-based learning due to the modeling errors and insufficient information in the offline dataset. In this section, we describe the implementation of SELFI on top of SACSoN.
4.1 Offline Learning with SACSoN